Font El Paragraph Paragraph Font cal styles 6. Given the following, find the DWL. P = 120 - 10Q MC = 20 FC = 0
Answer #1

P = 120 - 10Q

TR = P * Q

TR = (120 - 10Q) * Q

TR = 120Q - 10Q2

MR = STr/sQ = 120 - 20Q

Profit maximising level of output is produced at a point where the marginal revenue equals the marginal cost.

120 - 20Q = 20

120 - 20 = 20Q

100 = 20Q

Q = 100/20 = 5 units.

P = 120 - 10 * 5

P = $70.

Price DWT MC QEY 5 10

Area of yellow shaded triangle = 1/2 * base * height

DWL = 1/2 * (10 - 5) * (70 - 20)

DWL = $125.